# ColorBench: Can VLMs See and Understand the Colorful World? A Comprehensive Benchmark for Color Perception, Reasoning, and Robustness

### v2

@@ 4.1 VLM Benchmarks
4.2 Color Evaluation
Color understanding is increasingly recognized as a crucial aspect of Vision-Language Models’ ability to perceive and interpret visual content. Limited studies have explored how color information influences model performance on specific tasks. Some studies [33, 34] explore the understanding of color by replacing color-related words in textual inputs to evaluate the models’ ability to handle color-specific information. More recent research [35, 36] focus on assessing fine-grained color discrimination by asking models to distinguish subtle color differences in visual inputs. Samin et al. [37] introduced color-related foils to test VLMs’ capacity to cognize about basic colors like red, white, and green, particularly in contexts requiring attention to subtle cues. Additionally, Burapacheep et al. [38] developed a benchmark dataset to evaluate and enhance compositional color comprehension in VLMs, emphasizing tasks where understanding minimal color relationships is essential. IllusionVQA [39] evaluates model perception of color illusions in photorealistic scenes. While these works have addressed isolated aspects of color understanding, none has provided a holistic assessment framework. In contrast to these previous works, our study establishes the first comprehensive and specialized benchmark for evaluating the color-related abilities of VLMs, offering a quantitative, automated approach to further this area of research.
5 Conclusion, Limitation and Future Works
In this paper, we introduce ColorBench, the first benchmark designed to comprehensively evaluate the color understanding capabilities of VLMs, including Perception, Reasoning, and Robustness. After evaluating $32$ widely used VLMs on our benchmark, several undiscovered observations are revealed by us. These observations emphasize the need for more sophisticated model architectures that integrate deeper color reasoning capabilities. We plan to expand to more diverse tasks that involve complex interplays of color and texture, shape, or spatial relationships. Additionally, investigating the impact of different visual encoders and language models could further elucidate the pathways through which VLMs process color information.
Appendix A Data Sources
@@
Appendix B Detailed Generation Process for Robustness
For the Color Robustness, we evaluate the consistency of VLMs when faced with instances that differ only in the color of the visual input. To systematically assess this effect, we define $3$ recoloring strategies that determine which part of the image is altered: i) Target Segment, ii) Largest Segment, and iii) Entire Image. As mentioned in Table 6, Target Segment strategy recolors only the segment containing the object referenced in the question. This strategy ensures that the modification directly affects the model’s perception of task-relevant content. Largest Segment strategy alters the color of the largest segment that is irrelevant to the question, testing whether models are distracted by dominant but unrelated visual changes. In contrast, Entire Image strategy applies a global color shift to evaluate the model’s sensitivity to overall color variations. As summarized in Table 6, the first two strategies introduce localized modifications, while the third assesses robustness to broader image-wide color changes. By incorporating both task-relevant and irrelevant edits, our benchmark provides a comprehensive evaluation of VLMs’ ability to handle color perturbations across different contexts.
While generating color variations, we derive seed images from CV-Bench [16], a publicly available benchmark. For each seed image, as shown in Figure 3, we first employ a Grounded Segmentation Model (GAM) [17] to extract segments and their corresponding labels. We then apply the predefined recoloring strategies to determine the editing region. Once the editing region is determined, we modify the color of the corresponding region. In HSV color space, since Saturation and Value control the purity or brightness of the color, and only Hue controls the color of the part, we only adjust the Hue value in the HSV color space. Specifically, we shift the Hue by $90$°, $180$°, and $270$°. These three values ensure that the color manipulations cover significant perceptual differences across the color spectrum. These variations allow us to examine whether VLMs rely excessively on color information and whether they maintain consistency in their predictions despite substantial color shifts. This process produces nine variations per seed image, covering different strategies and degrees of color change to enable a comprehensive robustness assessment. To ensure interpretability, human experts filter out unnatural or negligible modifications, resulting in a final selection of 493 seed images for robustness evaluation.
@@ Appendix B Detailed Generation Process for Robustness
Appendix C ColorBench Categories and Questions
Table 7 provides a detailed description of each task, alongside representative figures and sample questions that effectively demonstrate the specific capabilities being tested. Cases are provided for each task in Figure 7 to 17.
@@ Appendix C ColorBench Categories and Questions
Appendix D Implementation Details
@@
Appendix E Evaluation Prompts
@@
You’ll be given an image, an instruction and some options. You have to select the correct one. Do not explain your reasoning. Answer with only the letter that corresponds to the correct option. Do not repeat the entire answer.
@@
You’ll be given an image, an instruction and some options. You have to select the correct one. Think step by step before answering. Then conclude with the letter that corresponds to the correct option. Make sure the option letter is in the parentheses like (X). Do not include ( or ) in the response except for the answer.
Appendix F Human Evaluation
@@
Appendix G Reasoning Models with Thinking Process
To comprehensively assess the performance of VLMs with thinking process on ColorBench, except for proprietary models with chain-of-thought(CoT) prompt, we additionally conduct experiments with GPT-o3 on perception and reasoning tasks. GPT-o3 is the most recent powerful proprietary VLMs that is trained to think before answer with reinforcement learning. We use API-version of GPT-o3 (2025-04-16) for evaluation. The result is shown in Table 8, together with results of CoT prompting and human evaluation.
The results presented in Table 8 indicate that human evaluators achieve the highest performance across the majority of tasks, except for three specific categories: Object Recognition (O’Recog), Color Proportion (C’Prop), and Color Comparison (C’Comp), where GPT-o3 holds the highest scores. The performance differences between GPT-o3 and human evaluators on O’Recog and C’Comp tasks are relatively minor (less than $3\%$). However, GPT-o3 substantially outperforms both humans and other VLMs on the C’Prop task, with an advantage exceeding $12\%$. This significant gap on C’Prop aligns with expectations, as humans generally exhibit lower sensitivity to precise quantitative measures. Meanwhile, GPT-o3 might benefit from including the capability to utilize analytical tools for precise image assessments and proportion estimations.
On the remaining tasks, GPT-o3 consistently outperforms GPT-4o (CoT) and Gemini-2-flash (CoT), except for the Color Blindness (C’Blind) task, where GPT-o3 trails GPT-4o (CoT) by $3.7\%$. The C’Blind task requires VLMs to accurately identify numbers or strings in an image which is composed of colored dots. This task demands capabilities of precise color recognition combined with a holistic spatial perception. One plausible reason for GPT-o3’s inferior performance is its longer and more complex reasoning path, which may lead to overthinking. This might cause the model to focus too much on local details or choices of tool, at the expense of the global and intuitive perception needed for this task.
@@ Appendix G Reasoning Models with Thinking Process
Appendix H More Visualizations
H.1 VLM Size & Model Performance for Each Task
@@
Figure 19 to 27 present detailed correlations between the log-scaled sizes of VLM parameters and the performance metrics for each task of Perception and Reasoning Categories. Deeper color represents higher accuracy. Each line represents a model family with the sizes growing from small to large. This visualization clearly shows the correlation between performances and model sizes, larger model leads to higher performance.
@@
H.2 Vision Size & Model Performance for Each Task
Figure 31 to 32 show detailed correlations between the log-scaled sizes of vision encoders and the performance metrics for each task of Perception and Reasoning Categories. Colors represent different model families. Models that have the same vision encoder sizes but with different LLM sizes are plotted as different points. Given that the majority of Vision-Language Models (VLMs) utilize a singular type of vision encoder, and that the sizes of these encoders generally range between 300-400M, it becomes challenging to assess the scaling effects within vision encoders.
@@ H.2 Vision Size & Model Performance for Each Task
H.3 Performance for Each Model Family on Each Task
Figures 33 to 39 illustrate task performance across different models within the same model families. In general, models with more parameters tend to perform better on the majority of tasks.
@@ H.3 Performance for Each Model Family on Each Task
Appendix I Samples Cases
I.1 Effect of CoT
@@
However, for Color Recognition and Object Recognition tasks, the improvement of involving slow thinking is limited, as these two tasks heavily rely on the accurate cognition of the vision encoder. The sample cases are shown in Figure 51 and 52. For Color Illusion task, adding reasoning steps causes the model to focus more on the misleading environment and the relationship between the environment and the foreground objects. This thinking negatively influences the model performance. A sample case is shown by Figure 53.
@@
I.2 Effect of Grayscale
For most of the tasks in ColorBench, colors are critical clues for VLMs to generate the answers. We highlight these cases in Figure 54 to 61.
However, for Color Illusion and Color Mimicry tasks, color clues might mislead VLMs to wrong answers, as shown in Figure 62 and 63.
@@ I.2 Effect of Grayscale
I.3 Failure with LLM and Vision

### v3

@@ A.1 VLM Benchmarks
A.2 Color Evaluation
Color understanding is increasingly recognized as a crucial aspect of Vision-Language Models’ ability to perceive and interpret visual content. Limited studies have explored how color information influences model performance on specific tasks. Some studies [51, 50] explore the understanding of color by replacing color-related words in textual inputs to evaluate the models’ ability to handle color-specific information. More recent research [16, 21] focuses on assessing fine-grained color discrimination by asking models to distinguish subtle color differences in visual inputs. Samin et al. [39] introduced color-related foils to test VLMs’ capacity to cognize basic colors like red, white, and green, particularly in contexts requiring attention to subtle cues. Additionally, Burapacheep et al. [3] developed a benchmark dataset to evaluate and enhance compositional color comprehension in VLMs, emphasizing tasks where understanding minimal color relationships is essential. IllusionVQA [40] evaluates model perception of color illusions in photorealistic scenes. While these works have addressed isolated aspects of color understanding, none have provided a holistic assessment framework. In contrast to these previous works, our study establishes the first comprehensive and specialized benchmark for evaluating the color-related abilities of VLMs, offering a quantitative, automated approach to further this area of research.
Appendix B Data Sources
@@
Appendix C Detailed Generation Process for Robustness
For the Color Robustness, we evaluate the consistency of VLMs when faced with instances that differ only in the color of the visual input. To systematically assess this effect, we define $3$ recoloring strategies that determine which part of the image is altered: i) Target Segment, ii) Largest Segment, and iii) Entire Image. As mentioned in Table 6, Target Segment strategy recolors only the segment containing the object referenced in the question. This strategy ensures that the modification directly affects the model’s perception of task-relevant content. Largest Segment strategy alters the color of the largest segment that is irrelevant to the question, testing whether models are distracted by dominant but unrelated visual changes. In contrast, Entire Image strategy applies a global color shift to evaluate the model’s sensitivity to overall color variations. As summarized in Table 6, the first two strategies introduce localized modifications, while the third assesses robustness to broader image-wide color changes. Importantly, only color attributes are altered without modifying object shapes or contextual elements, which preserves the overall realism of the image. By incorporating both task-relevant and irrelevant edits, our benchmark provides a comprehensive evaluation of VLMs’ ability to handle color perturbations across different contexts.
While generating color variations, we derive seed images from CV-Bench [42], a publicly available benchmark. For each seed image, as shown in Figure 3, we first employ a Grounded Segmentation Model (GAM) [38] to extract segments and their corresponding labels. We then apply the predefined recoloring strategies to determine the editing region. Once the editing region is determined, we modify the color of the corresponding region. In HSV color space, since Saturation and Value control the purity or brightness of the color, and only Hue controls the color of the part, we only adjust the Hue value in the HSV color space. Specifically, we shift the Hue by 9090°, 180180°, and 270270°. These three values ensure that the color manipulations cover significant perceptual differences across the color spectrum. This process produces nine variations per seed image, covering different strategies and degrees of color change to enable a comprehensive robustness assessment. To ensure interpretability, human experts filter out unnatural or negligible modifications, resulting in a final selection of 493 seed images for robustness evaluation. Additionally, we select questions that are color-invariant, which means answers remain valid regardless of whether the recoloring appears fully natural. This design choice isolates color variation as the sole variable of interest and prevents confounding effects from semantic or contextual changes. Through these steps, we evaluate whether VLMs rely excessively on color information and whether they maintain consistency in their predictions despite substantial color shifts.
@@ Appendix C Detailed Generation Process for Robustness
Appendix D ColorBench Categories and Questions
Table 7 provides a detailed description of each task, alongside representative figures and sample questions that effectively demonstrate the specific capabilities being tested. Cases are provided for each task in Figure 6 to 16.
@@ Appendix D ColorBench Categories and Questions
Appendix E Implementation Details
@@
Appendix F Evaluation Prompts
@@
You’$l$ be given an image, an instruction and some options. You have to select the correct one. Do not explain your reasoning. Answer with only the letter that corresponds to the correct option. Do not repeat the entire answer.
@@
You’$l$ be given an image, an instruction and some options. You have to select the correct one. Think step by step before answering. Then conclude with the letter that corresponds to the correct option. Make sure the option letter is in the parentheses like (X). Do not include ( or ) in the response except for the answer.
Appendix G Human Evaluation
@@
Appendix H Reasoning Models with Thinking Process
To comprehensively assess the performance of VLMs with the thinking process on ColorBench, except for proprietary models with chain-of-thought(CoT) prompt, we additionally conduct experiments with GPT-o3 on perception and reasoning tasks. GPT-o3 is the most recent powerful proprietary VLMs that is trained to think before answering with reinforcement learning. We use the API version of GPT-o3 (2025-04-16) for evaluation. The result is shown in Table 8, together with results of CoT prompting and human evaluation.
The results presented in Table 8 indicate that human evaluators achieve the highest performance across the majority of tasks, except for three specific categories: Object Recognition (O’Recog), Color Proportion (C’Prop), and Color Comparison (C’Comp), where GPT-o3 holds the highest scores. The performance differences between GPT-o3 and human evaluators on O’Recog and C’Comp tasks are relatively minor (less than $3\%$). However, GPT-o3 substantially outperforms both humans and other VLMs on the C’Prop task, with an advantage exceeding $12\%$. This significant gap on C’Prop aligns with expectations, as humans generally exhibit lower sensitivity to precise quantitative measures. Meanwhile, GPT-o3 benefits from including the capability to utilize analytical tools for precise image assessments and continuous exhaustive visual search [26] to obtain better proportion estimations.
On the remaining tasks, GPT-o3 consistently outperforms GPT-4o (CoT) and Gemini-2-flash (CoT), except for the Color Blindness (C’Blind) task, where GPT-o3 trails GPT-4o (CoT) by $3.7\%$. The C’Blind task requires VLMs to accurately identify numbers or strings in an image that is composed of colored dots. This task demands capabilities of precise color recognition combined with a holistic spatial perception. One plausible reason for GPT-o3’s inferior performance is its longer and more complex reasoning path, which may lead to overthinking. This might cause the model to focus too much on local details or choices of tool, at the expense of the global and intuitive perception needed for this task.
@@ Appendix H Reasoning Models with Thinking Process
Appendix I Qualitative Analysis of Failure Cases
To gain deeper insights into VLM failures on color-related tasks, we conduct a detailed case analysis using Qwen2.5-VL-3B and 7B models on different tasks. Following the attention visualization methodology of Zhang et al. [49], we focus on instances where the 3B model fails but the 7B model succeeds, allowing a clearer examination of the underlying capability differences. The visualizations of attention maps are shown in Figure 17 to 25.
For Color Perception tasks, we analyze the Color Recognition and Object Recognition tasks (excluding Color Extraction, which contains single-color color images). Our preliminary findings show that only a small number of failures arise from incorrect object localization. In most cases, both models correctly attend to the relevant regions but still produce incorrect predictions. This indicates that VLMs cannot accurately interpret color information, rather than deficiencies in visual grounding for these basic perception tasks.
For Color Reasoning tasks, tasks such as Color Proportion, Color Comparison, Color Counting, and Color Illusion require integrating visual information across the entire image without a clear focus point. Attention maps show that both 3B and 7B models exhibit similar focus patterns but generate different answers, implying that the divergence mainly originates from the language reasoning component rather than the visual encoder. For tasks with explicit perception targets, including Object Counting, Color Mimicry, and Color Blindness, both models attend to the correct regions, yet the 3B model often fails to produce accurate predictions. These results reveal that current VLMs remain weak in color interpretability even when their attention is properly aligned.
Figure 17: Visualized Attention Maps for Color Recognition Tasks.
Figure 18: Visualized Attention Maps for Object Recognition Tasks.
Figure 19: Visualized Attention Maps for Color Proportion Tasks.
Figure 20: Visualized Attention Maps for Color Comparison Tasks.
Figure 21: Visualized Attention Maps for Color Counting Tasks.
Figure 22: Visualized Attention Maps for Object Counting Tasks.
Figure 23: Visualized Attention Maps for Color Illusion Tasks.
Figure 24: Visualized Attention Maps for Color Mimicry Tasks.
Figure 25: Visualized Attention Maps for Color Blindness Tasks.
Appendix J Effect of Different Modalities
To investigate the impact of color information, we compare model performance on RGB versus grayscale images, thereby isolating the role of color within the image modality. To further explore the contribution of the image modality, we also conduct experiments using textual input only (questions and answer choices), where the original input images are substituted with pure black images of identical dimensions.
Table 9: Average Accuracy (%) across three input settings (Text-only, Grayscale+Text, RGB+Text) on Color Perception and Reasoning tasks.
Color Perception Color Reasoning P & R
VLMs: << 7B
Text-only 29.2 30.6 31.6 29.6 35.3 24.5 20.6 35.5 41.7 23.4 29.3
Gray+Text 25.9 33.5 42.7 29.1 37.1 23.2 23.3 42.4 53.7 23.0 32.1
RGB+Text 55.3 35.7 63.6 37.3 42.4 22.5 26.1 37.5 50.6 25.0 37.4
VLMs: 7B - 8B
Text-only 23.7 35.4 32.3 20.6 29.7 18.4 19.3 36.7 36.9 21.1 26.7
Gray+Text 25.2 35.7 46.0 27.8 41.3 22.2 27.5 48.2 58.7 23.6 34.2
RGB+Text 60.4 42.4 73.0 41.8 49.1 22.7 32.7 41.5 50.0 23.4 41.1
VLMs: 10B - 30B
Text-only 26.9 33.6 32.8 25.0 34.7 26.5 22.3 38.2 40.0 18.9 28.9
Gray+Text 26.8 37.9 46.8 22.5 46.5 22.4 30.1 43.0 60.3 26.0 35.0
RGB+Text 68.4 41.5 79.7 43.0 51.3 25.3 34.4 33.8 55.4 26.6 43.2
VLMs: 30B - 70B
Text-only 28.9 36.5 31.8 16.3 29.0 15.4 16.3 42.7 33.6 15.9 25.6
Gray+Text 28.7 42.1 51.2 26.3 49.9 24.3 25.6 48.8 65.1 22.7 36.7
RGB+Text 73.4 48.8 81.6 49.5 55.2 24.7 37.3 36.1 61.1 25.5 46.2
VLMs: >> 70B
Text-only 26.0 47.4 35.7 20.9 36.9 21.6 24.0 35.8 33.9 21.8 29.8
Gray+Text 25.3 40.9 54.6 25.3 51.0 21.8 28.6 44.6 54.3 26.1 36.1
RGB+Text 73.4 54.7 82.5 45.6 62.4 26.7 39.6 33.9 53.9 29.6 47.6
Table 9 presents the average accuracy across models grouped by LLM size. The result demonstrates that removing the visual modality (text-only setting) leads to the lowest performance across the majority of tasks. The performance differences among the three input settings allow us to disentangle the impact of textual input, image context (excluding color), and color information itself.
Notably, in tasks such as Color Recognition and Object Recognition, the performance gap between text-only and grayscale experiments is relatively small, whereas both are significantly outperformed by the RGB input setting. This suggests that color cues play a substantially more important role than either contextual visual or textual information in these tasks.
Appendix K Fine-tuning Experiments on ColorBench
We conduct a series of fine-tuning experiments to investigate model adaptation on specialized color-centric tasks. These experiments leverage three synthetic datasets designed for Color Extraction, Color Illusion, and Color Blindness. Using our synthetic data generation pipeline, we curate dedicated training sets for this purpose, with sample counts summarized in Table 10.
Table 10: Number of synthetic samples generated for fine-tuning experiments.
Task Number of Samples
Color Extraction 2400
Color Illusion 2400
Color Blindness 2280
To systematically assess the influence of different model components, we perform a comprehensive ablation study on Qwen2.5-VL-3B and Qwen2.5-VL-7B with the following settings:
  • MLP only
  • Vision encoder only
  • MLP + Vision encoder (jointly)
  • LLM (LoRA) only
  • LLM (LoRA) + MLP
  • LLM (LoRA) + Vision encoder
  • LLM (LoRA) + MLP + Vision encoder (jointly)
For configurations involving the LLM, we adopt the LoRA approach to update a subset of its parameters, while the remaining modules are fully fine-tuned.
Table 11: Accuracy (%) of Qwen2.5-VL (3B and 7B) under different training strategies across ColorBench tasks. Bold numbers indicate the best results within each model group.
Trainable Modules Color Perception Color Reasoning P&R
Model LLM (LoRA) MLP Vision C’Recog C’Extract O’Recog C’Prop C’Comp C’Count O’Count C’Illu C’Mimic C’Blind Overall
Qwen2.5-3B 72.4 38.5 74.0 43.8 48.5 22.6 25.2 43.0 45.7 24.2 41.1
71.1 53.1 75.3 50.0 49.5 22.5 26.2 45.2 44.3 25.5 43.6
73.7 53.1 79.2 46.3 45.5 29.4 27.2 48.4 47.1 25.5 44.4
75.0 56.3 75.3 47.5 49.5 28.4 25.2 46.2 47.1 28.0 45.2
71.1 75.0 70.1 45.0 51.5 26.5 27.2 45.2 47.1 27.4 46.2
69.7 77.1 74.0 40.0 53.5 23.5 32.0 51.6 45.7 37.6 48.8
71.1 75.0 71.4 46.3 49.5 25.5 27.2 49.4 48.6 31.4 46.7
72.4 75.0 71.4 45.0 51.5 24.3 32.0 46.2 50.0 28.0 47.1
Qwen2.5-7B 76.3 49.0 84.4 47.5 52.5 19.6 34.0 44.1 55.7 28.7 46.2
72.4 42.7 84.4 42.5 59.4 20.6 29.1 45.2 47.1 28.7 45.2
77.6 59.4 81.8 47.5 56.4 25.5 29.1 51.6 50.0 35.6 51.2
78.9 61.5 80.5 41.3 55.4 20.6 29.1 47.3 48.6 30.1 47.7
75.0 78.1 83.1 51.3 60.4 21.6 35.0 52.7 54.3 35.6 52.4
72.4 82.3 83.1 51.3 57.4 19.6 30.1 51.6 52.9 33.1 51.2
75.0 83.3 83.1 45.0 56.4 15.7 30.1 53.8 54.3 33.1 51.5
77.6 82.3 83.1 50.0 55.5 23.3 31.1 52.7 55.7 33.1 51.7
The evaluation results with finetuned VLMs are shown in Table 11. Overall, models that include LoRA fine-tuning on the LLM component consistently outperform those without it, exhibiting a substantial improvement in overall accuracy. Importantly, the improvements are not confined to the directly targeted tasks (Color Extraction, Color Illusion, Color Blindness). These experiments show that fine-tuning the model on part of tasks also produces notable gains on some ancillary reasoning tasks, including Color Proportion, and Color Comparison.
However, the transfer of knowledge is not universally positive. Certain tasks demonstrated limited or even negative performance transfer, indicating that fine-tuning exclusively on specialized color objectives does not guarantee generalization across the full spectrum of color perception and reasoning. This finding underscores that while targeted training enhances specialized abilities, a balanced and robust performance profile necessitates the inclusion of more diverse data and training objectives.
Appendix L More Visualizations
L.1 VLM Size & Model Performance for Each Task
@@ Appendix I Qualitative Analysis of Failure Cases
Figure 27 to 35 present detailed correlations between the log-scaled sizes of VLM parameters and the performance metrics for each task of Perception and Reasoning Categories. Deeper color represents higher accuracy. Each line represents a model family with the sizes growing from small to large. This visualization clearly shows the correlation between performances and model sizes, larger model leads to higher performance.
@@
L.2 Vision Size & Model Performance for Each Task
Figure 39 to 40 show detailed correlations between the log-scaled sizes of vision encoders and the performance metrics for each task of Perception and Reasoning Categories. Colors represent different model families. Models that have the same vision encoder sizes but with different LLM sizes are plotted as different points. Given that the majority of Vision-Language Models (VLMs) utilize a singular type of vision encoder, and that the sizes of these encoders generally range between 300-400M, it becomes challenging to assess the scaling effects within vision encoders.
@@
L.3 Performance for Each Model Family on Each Task
Figures 41 to 47 illustrate task performance across different models within the same model families. In general, models with more parameters tend to perform better on the majority of tasks.
@@
Appendix M Samples Cases
M.1 Effect of CoT
@@
However, for Color Recognition and Object Recognition tasks, the improvement of involving slow thinking is limited, as these two tasks heavily rely on the accurate cognition of the vision encoder. The sample cases are shown in Figure 59 and 60. For Color Illusion task, adding reasoning steps causes the model to focus more on the misleading environment and the relationship between the environment and the foreground objects. This thinking negatively influences the model performance. A sample case is shown by Figure 61.
@@
M.2 Effect of Grayscale
For most of the tasks in ColorBench, colors are critical clues for VLMs to generate the answers. We highlight these cases in Figure 62 to 69.
However, for Color Illusion and Color Mimicry tasks, color clues might mislead VLMs to wrong answers, as shown in Figure 70 and 71.
@@
M.3 Failure with LLM and Vision